ETHANE ASYMMETRIC C-H STRETCHING VIBRATIONAL SPECTRA

\underline{MARIA VILLA}, Departamento de Quimica. Universidad Autonoma Metropolitana, Av. San Rafael Atlixco 186, Col Vicentina, Iztapala, Mexico D.F. 09340, MEXICO; MA. LUISA SENENT, Departamento de Astrofisica Molecular e Infrarroja, Instituto de Estructura de la Materia, C.S.I.C., Serrano 113B, Madrid 28006, SPAIN; and DAVID R. HIDALGO, Departamento de Quimica. Universidad Autonoma Metropolitana, Av. San Rafael Atlixco 186, Col Vicentina, Iztapala, Mexico D.F. 09340.

In this work we continue studying the ethane vibrational spectra, being considered as a non-rigid molecule. The Θ internal rotational angle and the symmetric and anti-symmetric CH stretching normal modes will be the variables under study. Using ab-initio calculations a three dimentional potential is obtained, and finally a three dimentional Far-infrared and infrared spectra will be presented. The results obtained will be compared with the ones reported in the literature.